Mayan kings and many school sports teams are named for the puma, cougar, or mountain lion Felis concolor, the best jumper among animals. It can jump to a height of 10.4 ft when leaving the ground at an angle of 47.5
Answer #1

H = vo^2*sin^2(theta)/(2*g)

vo = sqrt(2*g*H)/sin(theta)

= sqrt(2*9.8*10.4*0.3048)/sin(47.5)

= 10.69 m/s
Answer #2

here ,

speed = sqrt(2*32.17 * 10.4)/sin(47.5)

= 35.1 feet/s
Answer #3

Height = u2 sin2(47.5)/2g

10.4 feet = 3.2 m
3.2 = u2 (0.34566)
u2 = 1.766
u=1.329 m/s
Answer #4

Using equation v2-u2 = 2as

we have v = 0, a = -g = -9.8m/s2 s = 10.4ft = 3.17m u = usin\Theta

Therefore usin\Theta = sqrt(2*9.8*3.17)

u = 7.882/(sin 47.5) = 10.69m/s at 47.5 degrees with horizontal
Answer #5

The vertical velocity is zero at the highest point, where ft. If u is the vertical velocity at take off, then:
0 = u^2 - 2gh or u = sqrt (2gh) = ... ft/s

If V is the take off velocity (not speed), then:
u = V sin 44.2
V = u / sin 44.2 = ... ft/s


25.75 ft/s, 36.94 ft/s